OpenStudy (anonymous):

Checking my math with this one. frac{6}{x(x-2)} + frac{3}{x}

OpenStudy (anonymous):

\[\frac{6}{x(x-2)} + \frac{3}{x}\] like that?

OpenStudy (anonymous):

you can use \[x(x-2)\] as a common denominator. so you only have to multiply the second part

OpenStudy (anonymous):

\[\frac{6}{x(x-2)} + \frac{3}{x} \frac{6}{x(x-2)} + \frac{3(x-2)}{x(x-2)} \frac{6}{x(x-2)} + \frac{3x - 6}{x(x-2)}\]

OpenStudy (anonymous):

\[\frac{6}{x(x-2)} + \frac{3}{x}\frac{x-2}{x-2}\] \[=\frac{6+3(x-2)}{x(x-2)}\]

OpenStudy (anonymous):

\[\frac{3x}{x(x-2)}\]

OpenStudy (anonymous):

what damein said!

OpenStudy (anonymous):

Should be my final answer, right?

OpenStudy (anonymous):

oh you are damien! sorry

OpenStudy (anonymous):

Haha, yup. I was just wanting to check that I came to the right answer :)

OpenStudy (anonymous):

yup you got it

OpenStudy (anonymous):

Cool, thanks. And, is there a way to start the question off using the equation widget? That's the reason I didn't use it in my first post, because I couldn't get it to work.

OpenStudy (anonymous):

not sure. i just type in the latex, but you can always post and then type it in